OpenStudy (anonymous):

(MEDAL+FAN+COOKIE_HUG! YASSS) 8. Short Answer. A communication satellite transmits a radio wave at frequency of 5x10^9 Hz. What is the signal's wavelength? Assume the wave travels in a vacuum. Show your work. (3 points)

OpenStudy (ookawaiioo):

Give me a minute

OpenStudy (anonymous):

Oki ^_^

OpenStudy (ookawaiioo):

C (speed of light) = 3.0x10^8 m/s = wavelength * frequency. In waves problems, frequency is inversely proportional to wavelength. If the frequency of the radiowave (transverse wave) is 5.0x10^9 3.0x10^8 = 5.0x10^9 * wavelength (3.0x10^8/5.0x10^9) = Wavelength = 0.06 meters

OpenStudy (anonymous):

can I ask another question?

OpenStudy (ookawaiioo):

Go ahead.......

OpenStudy (anonymous):

19. Both gamma rays and X-rays are used to see inside the body. Which one is used to make images of bones? how are the other rays used?

OpenStudy (ookawaiioo):

X-ray is used to make images of bones. Bones are very dense with Calcium nuclei which makes them very visible on the x-ray films. Gamma ray = it would kill you. Not specifically should on the applications of gamma rays. Probably in nuclear chemistry, with nuclear fission and decay.

OpenStudy (anonymous):

wow, thank you! ^_^ you're really smart♥

OpenStudy (ookawaiioo):

Thanks, and goodluck on your endeavors.